OpenStudy (anonymous):

i have the answer for the problem but i dont know how to solve it... find a polynomial function with real coefficients and of least possible degree having the given zeros. Let the leading coefficient be 1 8,2,3

OpenStudy (anonymous):

f(x) = (x-8)(x-2)(x-3)

OpenStudy (anonymous):

am I just suppose to distribute??

OpenStudy (anonymous):

If a polynomial has those zeros, then: x=8 x=2 x=3 think of it like a quadratic so, x-8=0 x-2=0 x-3=0 therefore (x-8)(x-2)(x-3) = f(x)

OpenStudy (anonymous):

yeah, just distribute that

OpenStudy (anonymous):

Thanks!